# Defending Unauthorized Model Merging via Dual-Stage Weight Protection

### v1

@@ 1 Introduction
MergeGuard has shown high effectiveness on a variety of tasks and architectures. For ViT-L-14 on the GTSRB and MNIST, the accuracy of the preprocessed model only decreased slightly (less than 1%). However, after merging, the accuracy of GTSRB plummeted from 86.78% to 12.19%, and MNIST from 96.02% to 11.35%. For large language models (LLMs), the accuracy after merging generally dropped to single digits, with some even falling to 0%. For example, the accuracy of the Gemma2 decreased from 69.6% to 1.52% on the GSM8K after merging, and from 64.02% to 21.34% on the HumanEval.
@@
  • Experiments on ViT-L-14, Llama2, Gemma2, and Mistral show that MergeGuard reduces merged-model accuracy by up to 90% with less than 1.5% loss on the protected model.
@@ 4 System Model
Free-rider’s Capability. The free-rider is assumed to have limited computational resources, making expensive retraining or knowledge distillation [15, 36] infeasible. In particular, because the free-rider is highly cost-sensitive, they will not perform knowledge distillation on $T_{\mathrm{def}}$ even though the dataset is downloadable, as such retraining still incurs substantially higher computation and tuning costs compared to direct parameter merging. Instead, they rely on low-cost model merging as a shortcut to acquire the defender’s expertise. The free-rider has full white-box access to the defender’s released model and a task-specific model $\theta_{\mathrm{fr}}$ derived from the same $\theta_{\mathrm{pre}}$. The attack is considered successful if the merged model $\theta_{\mathrm{merge}}$ achieves accuracy comparable to $\theta_{\mathrm{def}}$ on the defender’s task $T_{\mathrm{def}}$ while preserving its own performance on $T_{\mathrm{fr}}$, where $T_{X}$ denotes $X$’s domain-specific task.
@@
Defender’s Capability. The defender has full access to and control over the parameters of their own model $\theta_{\mathrm{def}}$ and can apply parameter-level transformations to strengthen protection. These modifications must maintain the model’s original task performance while preventing the merged model from retaining it. Although the defender cannot access the free-rider’s model, it can reasonably assume both originate from the same model $\theta_{\mathrm{pre}}$; otherwise, merging would not be feasible [40, 16, 10, 6, 12, 24, 46]. A successful defense ensures that the released model remains fully functional for legitimate use but any merged version experiences a substantial accuracy drop on $T_{\mathrm{def}}$.
@@
Problem Setup. The free-rider’s goal is to construct a merged model $\theta_{\mathrm{merge}}=\text{Merge}(\theta_{\mathrm{pre}},\theta_{\mathrm{def}},\theta_{\mathrm{fr}})$ that simultaneously satisfies $\mathrm{Perf}(\theta_{\mathrm{merge}},T_{\mathrm{def}})\approx\mathrm{Perf}(\theta_{\mathrm{def}},T_{\mathrm{def}})$ and $\mathrm{Perf}(\theta_{\mathrm{merge}},T_{\mathrm{fr}})\approx\mathrm{Perf}(\theta_{\mathrm{fr}},T_{\mathrm{fr}})$, where $\mathrm{Merge}(\cdot)$ denotes a parameter-level fusion method such as model soups or task arithmetic [40, 16, 10, 6, 12, 24, 46], and $\mathrm{Perf}(\cdot)$ represents the appropriate utility metric for the target model family (e.g., accuracy for vision or perplexity for language models). To counter this, the defender instead releases a protected model $\hat{\theta}_{\mathrm{def}}=\mathrm{Defense}(\theta_{\mathrm{def}})$ whose standalone performance remains intact, $\mathrm{Perf}(\hat{\theta}_{\mathrm{def}},T_{\mathrm{def}})\approx\mathrm{Perf}(\theta_{\mathrm{def}},T_{\mathrm{def}})$, but any merged version $\hat{\theta}_{\mathrm{merge}}=\text{Merge}(\theta_{\mathrm{pre}},\hat{\theta}_{\mathrm{def}},\theta_{\mathrm{fr}})$ exhibits a pronounced degradation, $\mathrm{Perf}(\hat{\theta}_{\mathrm{merge}},T_{\mathrm{def}})\ll\mathrm{Perf}(\theta_{\mathrm{def}},T_{\mathrm{def}})$. Hence, our objective is to design a proactive defense function $\mathrm{Defense}(\cdot)$ that enforces this degradation property while maintaining the protected model’s utility on its intended task.
@@ 5 Proposed Method
[…]
@@
As discussed in Section 3, most of recent model merging techniques such as TIES [42] and AdaMerging [46] rely on sparsification to distribute task-relevant parameters sparsely across layers, reducing interference and enabling merged models to preserve multiple capabilities.
@@
The goal of this stage is to disperse task-critical weights across different layers through training. In addition to the standard cross-entropy loss $L_{\mathrm{CE}}$ used in classification, we introduce an $L_{2}$ regularization term that drives model parameters toward zero, encouraging a more balanced distribution of important information. The overall loss can be expressed as:
$L_{\mathrm{total}}=L_{\mathrm{CE}}+\alpha L_{2},$ (1)
where $\alpha$ is a hyperparameter that balances $L_{\mathrm{CE}}$ and $L_{2}$. For vision models, the loss takes the form $L_{\mathrm{total}}=-\sum_{t=1}^{T}y_{t}\log(\hat{y}_{t})+\alpha\sum_{\ell=1}^{L}\|\theta^{(\ell)}\|_{2}^{2}$, where $L_{\mathrm{CE}}=-\sum_{t=1}y_{t}\log(\hat{y}_{t})$ and $L_{2}=\sum_{\ell=1}^{L}\|\theta^{(\ell)}\|_{2}^{2}$. Here, $y_{t}$ denotes the ground-truth label of the $t$-th class, $\hat{y}_{t}$ the predicted probability, and $\theta^{(\ell)}$ the trainable parameters of layer $\ell$. The $L_{2}$ penalty is applied independently to each layer.
For LLMs, the loss is written as:
@@
where $P(y_{t}\mid y_{<t},x;\theta)$ is the conditional probability of predicting token $y_{t}$ given the previous context $y_{<t}$. This formulation applies $L_{2}$ regularization to each layer’s parameters, smoothing large-magnitude weights and evenly distributing task-relevant signals across the network.
@@
This stage introduces a training-free procedure that intentionally offsets task-relevant weights to disrupt merging compatibility. We begin by masking each layer of the model and measuring the resulting accuracy drop. The top $k^{\prime}\%$ of layers that cause the largest performance degradation are identified as critical and excluded from subsequent operations.
Afterwards, a binary masking vector $M$ is then constructed, where elements corresponding to $k^{\prime}\%$ critical layers and to the least significant $(1-k)(1-k^{\prime})\%$ parameters are set to zero, and those in the remaining $k(1-k^{\prime})\%$ range are set to one. This selective preservation ensures that later weight adjustments do not damage essential functionality.
Next, we calculate
@@
We particularly note that some layers differ drastically in parameter scale across tasks, and directly applying this operation can destabilize the model. To prevent such collapse in LLMs, we exclude three layers, model.embed_tokens.weight, model.norm.weight, and lm_head.weight, from the subtraction process, while ViT-based models undergo the full operation.
@@
MergeGuard breaks this assumption by jointly modifying both the magnitude distribution and the directional alignment of $\tau_{\mathrm{def}}$. Formally, after Stage 1, the finetuned parameters become $\theta^{\prime}_{\mathrm{def}}=\theta_{\mathrm{pre}}+\tau^{\prime}_{\mathrm{def}}$, where $\tau^{\prime}_{\mathrm{def}}=\tau_{\mathrm{def}}+\epsilon$, $\mathbb{E}[\epsilon]=0$, and $\text{Cov}[\epsilon]=\sigma^{2}I$. This step redistributes task-relevant gradients isotropically, flattening the curvature of the local basin and ensuring that important information is evenly spread rather than concentrated along a few dominant directions. Consequently, the Hessian eigenvectors associated with large eigenvalues, those that previously spanned the shared merging subspace, become decorrelated and dispersed.
@@ 6.1 Experiment Setups
Implementation of MergeGuard. In our implementation, the Density-Aware Finetuning stage requires no additional hyperparameter tuning. The Adversarial Weight Negation stage introduces two hyperparameters, $k^{\prime}$ and $k$, controlling the proportion of preserved and perturbed layers, respectively. Unless otherwise stated, we set $k^{\prime}=10$ and $k=0.1$ throughout all experiments. $\alpha$ in Eq. 1 is 0.01 and $\beta$ in Stage 2 is set to 1 in our experiments. The merging coefficients are determined by each model merging methods [40, 16, 42, 46, 48].
@@ 6.2.1 Results on Image Classification
Table 3 summarizes the image classification results comparing MergeGuard with PaRaMS [18]. During the preprocessing stage, MergeGuard introduces a minor accuracy drop on the original model due to the regularization and perturbation applied in the protection process. In contrast, PaRaMS [18], which only performs parameter rearrangement, preserves the original accuracy almost perfectly. However, after model merging, the gap between the two defenses becomes substantial. MergeGuard consistently enforces stronger degradation on the merged models while maintaining near-original accuracy on the protected model. For instance, on GTSRB and MNIST, MergeGuard in TA reduced the original accuracy by less than 2%, yet the merged accuracy dropped sharply from 86.78% to 12.91% and from 98.94% to 11.35%, respectively. By comparison, PaRaMS only reduced the same metrics to 52.25% and 96.02%. These results confirm that MergeGuard substantially weakens merging compatibility.
@@
Table 4: The results of ablation study. Each number indicates classification accuracy (%).
@@ 6.2.2 Results on Image Generation
Figure 3 illustrates the qualitative comparison on Stable Diffusion 1.5. We use Naruto as the defender’s task ($T_{\mathrm{def}}$) and WikiArt as the free-rider’s task ($T_{\mathrm{fr}}$). The first two columns visualize the images generated by $\theta_{\mathrm{def}}$ and $\theta_{\mathrm{fr}}$, respectively. When tested with prompts $A_{0}$ and $A_{1}$ designed for Naruto, the free-rider’s model $\theta_{\mathrm{fr}}$ can also reproduce the character, indicating that it has captured the semantic style of Naruto. The merged model $\theta_{\mathrm{merge}}$, obtained through standard parameter fusion, inherits this capability and successfully generates Naruto-style images, demonstrating a clear case of intellectual property leakage.
In contrast, under MergeGuard, the protected standalone model $\hat{\theta}_{\mathrm{def}}$ still preserves its generative fidelity and produces faithful Naruto imagery. However, the protected merged model $\hat{\theta}_{\mathrm{merge}}$ fails to reproduce the Naruto character, yielding only partially coherent but semantically mismatched results. This degradation confirms that MergeGuard effectively prevents the unauthorized inheritance of generative capabilities while preserving the original model’s usability.
@@ 6.2.3 Results on Text Classification
Figure 4 presents the performance of MergeGuard on LLMs under various merging strategies. Across all benchmarks, the protected models maintain comparable accuracy prior to merging. However, once merged, their performance degrades sharply. For instance, both Llama2 and Mistral experience nearly a 50% drop in overall accuracy, while Gemma2 suffers even more severe degradation. On the GSM8K benchmark, which evaluates mathematical reasoning, the merged models become almost nonfunctional, showing frequent calculation errors and skipped reasoning steps, with accuracy often falling to single-digit levels (e.g., Mistral with TA drops from 75.7% to 0%). Similarly, on the HumanEval code generation task, MergeGuard consistently reduces post-merge pass@1 accuracy by more than half. Qualitative inspection further reveals that many generated programs, though sometimes passing test cases, contain redundant or looping code segments, as illustrated in the Appendix.
@@ 6.4 Discussion
This discrepancy stems from how each method disrupts representational continuity in generative models. Diffusion-based generators depend on smooth manifold alignment between latent and pixel spaces. PaRaMS introduces random multi-head scaling and parameter rearrangement, which destroy such continuity, leading to decorrelated feature channels and complete breakdown of generative structure. In contrast, MergeGuard perturbs model parameters in a structured manner: its $L_{2}$-regularized redistribution preserves weak weight correlations that maintain partial coherence in the latent manifold. Thus, MergeGuard prevents semantic fidelity without inducing total noise collapse.
@@ 7 Adaptive Attack
Summary of Adaptive Attacks. For Unmask, although subtracting the protected model’s parameters can partially neutralize MergeGuard’s perturbations, it concurrently removes essential task-specific information, causing the merged model’s performance on $T_{\mathrm{def}}$ to regress to the pretrained level $T_{\mathrm{pre}}$, as shown in Table 5. In contrast, GradErase projects each gradient update orthogonally to the estimated disturbance vector $v_{\mathrm{disturb}}$, thereby avoiding reinforcement of the defended direction. However, because the estimation of $v_{\mathrm{disturb}}$ is inherently noisy and incomplete, this attack fails to fully restore merging compatibility.
Overall, as summarized in Table 5, MergeGuard reduces the merged accuracy from $84.47$ to $40.26$ on average. Although Unmask and GradErase can partially recover the merged model to $64.64$ and $51.77$, respectively, MergeGuard still maintains a substantial degradation margin of at least $20\%$ across all tasks. This demonstrates that even under adaptive settings, MergeGuard remains effective in preventing the merged model from achieving functional usability.

### v2

@@ 1 Introduction
MergeGuard has shown high effectiveness on a variety of tasks and architectures. For ViT-L-14 on the GTSRB and MNIST, the accuracy of the preprocessed model only decreased slightly (less than 11%). However, after merging, the accuracy of GTSRB plummeted from 86.7886.78% to 12.1912.19%, and MNIST from 96.0296.02% to 11.3511.35%. For large language models (LLMs), the accuracy after merging generally dropped to single digits, with some even falling to 0%. For example, the accuracy of the Gemma2 decreased from 69.669.6% to 1.521.52% on the GSM8K after merging, and from 64.0264.02% to 21.3421.34% on the HumanEval.
@@
  • Experiments on ViT-L-14, Llama2, Gemma2, and Mistral show that MergeGuard reduces merged-model accuracy by up to 9090% with less than 1.51.5% loss on the protected model.
@@ 4 System Model
Free-rider’s Capability. The free-rider is assumed to have limited computational resources, making expensive retraining or knowledge distillation [15, 36] infeasible. In particular, because the free-rider is highly cost-sensitive, they will not perform knowledge distillation on $T_{\mathrm{def}}$ even though the dataset is downloadable, as such retraining still incurs substantially higher computation and tuning costs than direct parameter merging. Instead, they rely on low-cost model merging as a shortcut to acquire the defender’s expertise. The free-rider has full white-box access to the defender’s released model and a task-specific model $\theta_{\mathrm{fr}}$ derived from the same $\theta_{\mathrm{pre}}$. The attack is considered successful if the merged model $\theta_{\mathrm{merge}}$ achieves accuracy comparable to $\theta_{\mathrm{def}}$ on the defender’s task $T_{\mathrm{def}}$ while preserving its own performance on $T_{\mathrm{fr}}$, where $T_{X}$ denotes $X$’s domain-specific task.
@@
Defender’s Capability. The defender has full access to and control over the parameters of their own model $\theta_{\mathrm{def}}$ and can apply parameter-level transformations to strengthen protection. These modifications must maintain the model’s original task performance while preventing the merged model from retaining it. Although the defender cannot access the free-rider’s model, it can reasonably assume both originate from the same model $\theta_{\mathrm{pre}}$; otherwise, merging would not be feasible [40, 16, 10, 6, 12, 24, 46]. A successful defense ensures that the released model remains fully functional for legitimate use, but any merged version experiences a substantial accuracy drop on $T_{\mathrm{def}}$.
@@
Problem Setup. The free-rider’s goal is to construct a merged model $\theta_{\mathrm{merge}}=\text{Merge}(\theta_{\mathrm{pre}},\theta_{\mathrm{def}},\theta_{\mathrm{fr}})$ that simultaneously satisfies $\mathrm{Perf}(\theta_{\mathrm{merge}},T_{\mathrm{def}})\approx\mathrm{Perf}(\theta_{\mathrm{def}},T_{\mathrm{def}})$ and $\mathrm{Perf}(\theta_{\mathrm{merge}},T_{\mathrm{fr}})\approx\mathrm{Perf}(\theta_{\mathrm{fr}},T_{\mathrm{fr}})$, where $\mathrm{Merge}(\cdot)$ denotes a parameter-level fusion method such as model soups or task arithmetic [40, 16, 10, 6, 12, 24, 46], and $\mathrm{Perf}(\cdot)$ represents the appropriate utility metric for the target model family (e.g., accuracy for vision or perplexity for language models). To counter this, the defender instead releases a protected model $\hat{\theta}_{\mathrm{def}}=\mathrm{Defense}(\theta_{\mathrm{def}})$ whose standalone performance remains intact, $\mathrm{Perf}(\hat{\theta}_{\mathrm{def}},T_{\mathrm{def}})\approx\mathrm{Perf}(\theta_{\mathrm{def}},T_{\mathrm{def}})$, but any merged version $\hat{\theta}_{\mathrm{merge}}=\text{Merge}(\theta_{\mathrm{pre}},\hat{\theta}_{\mathrm{def}},\theta_{\mathrm{fr}})$ exhibits a pronounced degradation, $\mathrm{Perf}(\hat{\theta}_{\mathrm{merge}},T_{\mathrm{def}})\ll\mathrm{Perf}(\theta_{\mathrm{def}},T_{\mathrm{def}})$. Therefore, our objective is to design a proactive defense function $\mathrm{Defense}(\cdot)$ that enforces this degradation property while maintaining the protected model’s utility on its intended task.
@@ 5 Proposed Method
To address unauthorized merging, we propose MergeGuard, which comprises two stages: 1 Density-Aware Finetuning (training stage) and 2 Adversarial Weight Negation (training-free stage). We first outline the key idea behind MergeGuard, then describe each stage in detail.
@@
As discussed in Section 3, most recent model merging techniques, such as TIES [42] and AdaMerging [46] rely on sparsification to distribute task-relevant parameters sparsely across layers, reducing interference and enabling merged models to preserve multiple capabilities.
@@
The goal of this stage is to disperse task-critical weights across different layers through training. In addition to the standard cross-entropy loss $L_{\mathrm{CE}}$ used in classification, we introduce an $L_{2}$ regularization term that drives model parameters toward zero, encouraging a more balanced distribution of important information. The total loss function is formulated as:
$L_{\mathrm{Total}}=L_{\mathrm{CE}}+\alpha L_{2}$ (1)
Specifically, for vision tasks, the loss can be explicitly expanded as:
$L_{\mathrm{Total}}=-\sum_{t=1}^{T}y_{t}\log(\hat{y}_{t})+\alpha\sum_{\ell=1}^{L}\|\theta^{(\ell)}\|_{2}^{2},$ (2)
where $y_{t}$ and $\hat{y}_{t}$ denote the ground-truth label and the predicted probability for the $t$-th class, respectively; $\alpha$ is a hyperparameter balancing the two terms; and $\theta^{(\ell)}$ represents the trainable parameters of the $\ell$-th layer. Crucially, the $L_{2}$ penalty is applied layer-wise so that each layer is regularized independently.
Similarly, for LLMs, the objective function is formulated based on next-token prediction:
@@
where $P(y_{t}\mid y_{<t},x;\theta)$ denotes the conditional probability of predicting the token $y_{t}$ given the preceding context $y_{<t}$ and input $x$. This formulation applies $L_{2}$ regularization across all layer parameters, effectively smoothing large-magnitude weights and promoting an even distribution of task-relevant signals throughout the network layers.
@@
This stage introduces a training-free procedure that intentionally offsets task-relevant weights to disrupt merging compatibility. We begin by masking each layer of the model and measuring the resulting accuracy drop. The top $k^{\prime}\%$ of layers that cause the largest performance degradation are identified as critical and excluded from subsequent operations. Afterwards, a binary masking vector $M$ is then constructed, where elements corresponding to $k^{\prime}\%$ critical layers and to the least significant $(1-k)(1-k^{\prime})\%$ parameters are set to zero, and those in the remaining $k(1-k^{\prime})\%$ range are set to one. This selective preservation ensures that later weight adjustments do not damage essential functionality. Next, we calculate
@@
We note that some layers differ substantially in parameter scale across tasks, and applying this operation directly can destabilize the model. To prevent collapse in LLMs, we exclude three layers, model.embed_tokens.weight, model.norm.weight, and lm_head.weight, from the subtraction process, while ViT-based models undergo the full operation.
@@
MergeGuard breaks this assumption by jointly modifying both the magnitude distribution and the directional alignment of $\tau_{\mathrm{def}}$. Formally, after Stage 1, the finetuned parameters become $\theta^{\prime}_{\mathrm{def}}=\theta_{\mathrm{pre}}+\tau^{\prime}_{\mathrm{def}}$, where $\tau^{\prime}_{\mathrm{def}}=\tau_{\mathrm{def}}+\epsilon$, $\mathbb{E}[\epsilon]=0$, and $\text{Cov}[\epsilon]=\sigma^{2}I$. This step redistributes task-relevant gradients more isotropically, flattening the curvature of the local basin and ensuring that important information is evenly spread rather than concentrated along a few dominant directions. As a result, the Hessian eigenvectors associated with large eigenvalues that previously spanned the shared merging subspace become decorrelated and dispersed.
@@ 6.1 Experiment Setups
Implementation of MergeGuard. Our implementation of the Density-Aware Finetuning stage requires no additional hyperparameter tuning. The Adversarial Weight Negation stage introduces two key hyperparameters: $k^{\prime}$ and $k$, which control the proportion of preserved and perturbed layers, respectively. Unless otherwise specified, we set $k^{\prime}=10$ and $k=0.1$ for all experiments. In Eq. 1, we set $\alpha=0.01$, while $\beta$ in Stage 2 is set to $1$. Merging coefficients are determined following the protocols of the respective model merging methods [40, 16, 42, 46, 48].
@@ 6.2.1 Results on Image Classification
Table 3 summarizes the image classification results comparing MergeGuard with PaRaMS [18]. During the preprocessing stage, MergeGuard introduces a minor accuracy drop on the original model due to the regularization and perturbation applied in the protection process. In contrast, PaRaMS [18], which only performs parameter rearrangement, preserves the original accuracy almost perfectly. However, after model merging, the gap between the two defenses becomes substantial. MergeGuard consistently enforces stronger degradation on the merged models while maintaining near-original accuracy on the protected model. For instance, on GTSRB and MNIST, MergeGuard in TA reduced the original accuracy by less than 22%, yet the merged accuracy dropped sharply from 86.7886.78% to 12.9112.91% and from 98.9498.94% to 11.3511.35%, respectively. By comparison, PaRaMS only reduced the same metrics to 52.2552.25% and 96.0296.02%. These results confirm that MergeGuard substantially weakens merging compatibility.
@@
Table 4: The results of the ablation study. Each number indicates classification accuracy (%).
@@ 6.2.2 Results on Image Generation
Figure 3 illustrates the qualitative comparison on SD 1.5. We use Naruto as the defender’s task ($T_{\mathrm{def}}$) and WikiArt as the free-rider’s task ($T_{\mathrm{fr}}$). The first two columns visualize the images generated by $\theta_{\mathrm{def}}$ and $\theta_{\mathrm{fr}}$, respectively. When tested with prompts $A_{0}$ and $A_{1}$ designed for Naruto, the free-rider’s model $\theta_{\mathrm{fr}}$ can also reproduce the character, indicating that it has captured the semantic style of Naruto. The merged model $\theta_{\mathrm{merge}}$, obtained through standard parameter fusion, inherits this capability and successfully generates Naruto-style images, demonstrating a clear case of intellectual property leakage.
In contrast, under MergeGuard, the protected standalone model $\hat{\theta}_{\mathrm{def}}$ maintains its generative fidelity, producing faithful Naruto imagery. However, the protected merged model $\hat{\theta}_{\mathrm{merge}}$ fails to replicate the character, yielding only incoherent or semantically mismatched outputs. This degradation confirms that MergeGuard effectively prevents the unauthorized inheritance of generative capabilities without compromising the original model’s utility.
@@ 6.2.3 Results on Text Classification
Figure 4 illustrates MergeGuard’s performance across various LLM merging strategies. While protected models maintain competitive accuracy prior to merging, their performance collapses post-merge across all benchmarks. For instance, both Llama2 and Mistral experience nearly a 5050% drop in overall accuracy, while Gemma2 suffers even more severe degradation. Notably, on the GSM8K mathematical reasoning benchmark, merged models become virtually nonfunctional, frequently skipping reasoning steps or making elementary calculation errors, with accuracy often plunging to single digits (e.g., Mistral with TA drops from 75.775.7% to 0%). Similarly, on the HumanEval code generation task, MergeGuard consistently reduces post-merge pass@1 accuracy by over 50%. Qualitative analysis (see Appendix) further reveals that while some generated programs pass basic test cases, they often contain redundant logic or infinite loops.
@@ 6.4 Discussion
This discrepancy stems from how each method disrupts representational continuity in generative models. Diffusion-based generators depend on smooth manifold alignment between latent and pixel spaces. PaRaMS introduces random multi-head scaling and parameter rearrangement, which destroy such continuity, leading to decorrelated feature channels and a complete breakdown of the generative structure. In contrast, MergeGuard perturbs model parameters in a structured manner: its $L_{2}$-regularized redistribution preserves weak weight correlations that maintain partial coherence in the latent manifold. Thus, MergeGuard prevents semantic fidelity without inducing total noise collapse.
@@ 7 Adaptive Attack
Summary of Adaptive Attacks. For Unmask, although subtracting the protected model’s parameters can partially neutralize MergeGuard’s perturbations, it concurrently removes essential task-specific information, causing the merged model’s performance on $T_{\mathrm{def}}$ to regress to the pretrained level $T_{\mathrm{pre}}$, as shown in Table 5. In contrast, GradErase projects each gradient update orthogonally to the estimated disturbance vector $v_{\mathrm{disturb}}$ to avoid reinforcing the defended direction. However, as the estimation of $v_{\mathrm{disturb}}$ is inherently noisy and incomplete, this attack fails to fully restore merging compatibility.
As summarized in Table 5, MergeGuard reduces the average merged accuracy from $84.47\%$ to $40.26\%$. Although adaptive attacks such as Unmask and GradErase can partially recover performance to $64.64\%$ and $51.77\%$ respectively, MergeGuard consistently maintains a substantial degradation margin of at least $20\%$ across all tasks. These results demonstrate that even under adversarial settings, MergeGuard remains robust, effectively preventing the merged model from regaining functional usability.